# Effects of Artificial Collisions, Filtering, and Nonlocal Closure Approaches on Hermite-based Vlasov-Poisson Simulations

### v1

@@ 1 Introduction
Kinetic simulations of collisionless plasmas are important for many applications, e.g. astrophysics, geophysics, and laboratory fusion devices. Solving the collisionless Vlasov equation numerically is challenging due to its high dimensionality, Hamiltonian structure, and significant spatial and temporal scale separation. Furthermore, kinetic solvers struggle to accurately capture phase-space mixing, in which the distribution function develops increasingly fine structures in velocity space, known as filamentation [1]. As shown by the solution $f(x,v,t)=f(x-vt,v,0)$ to the free-streaming equation $\partial_{t}f+v\partial_{x}f=0$, these structures cannot be resolved after a finite time determined by the resolution of the velocity space. Filamentation can also cause recurrence effects in Eulerian solvers (grid or spectral), where the solution is artificially periodic in time [2, 3].
@@
The most common closure for the Hermite spectral approximation is closure by truncation, which assumes that the last spectral moment is zero. However, the Landau root [26] does not emerge as a discrete eigenvalue with the closure by truncation, regardless of velocity resolution. Grant and Feix [1] were the first to show that adding artificial collisions to the finite system with the closure by truncation can restore the correct Landau damping [27, 28, 20]. Previous work by [1, 9] use the Lenard-Bernstein (LB) [29] collisional operator. Similarly, recent work by [23, 15, 16, 11, 12, 13, 17] use a hypercollisional operator, which damps more strongly the higher-order Hermite moments in comparison to the LB operator. Filtering techniques, such as Klimas [30] and Hou-Li [31], are also used to suppress recurrence and filamentation artifacts [32, 33, 34, 35, 3]. Another method for incorporating kinetic physics into moment solvers is with a linear nonlocal closure, often referred to as Landau-fluid closures [36, 37]. This approach matches the closure coefficients to the asymptotic limits of the linear kinetic response function. Landau fluid closures were initially formulated for fluid models incorporating three- and four-moments and were later generalized to models involving higher moments in Hermite space [38]. Therefore, it becomes essential to incorporate a modified term or closure that introduces dissipation into the discrete system to prevent filamentation issues while ensuring that the chosen approach accurately captures the system’s average dynamics.
@@ 2.2 Linear Response Function
$-i\omega\hat{f}(v)+ikv\hat{f}(v)+ik\hat{\phi}\frac{\mathrm{d}f_{0}}{\mathrm{d}v}=0\qquad\Longrightarrow\qquad\hat{f}(v)=k\hat{\phi}\frac{\mathrm{d}f_{0}/\mathrm{d}v}{\omega-kv}.$ (4)
For a Maxwellian equilibrium, that is, $f_{0}(v)=\exp\left(-v^{2}/2\right)/\sqrt{2\pi}$, the perturbed density is described via the linear kinetic response function $R(\xi)$, i.e.,
@@
where the plasma dispersion function is $Z(\xi)=\pi^{-1/2}\int_{\mathbb{R}}\exp(-s^{2})/(s-\xi)\mathrm{d}s$ for $\mathrm{Im}(\xi)>0$ and is analytically continued for $\mathrm{Im}(\xi)\leq 0$ [42]. The asymptotic expansion of the response function is given by [40, §3]:
@@ 2.3 Asymmetrically Weighted Hermite Expansion
We discretize the electron distribution function in velocity space via an asymmetrically weighted (AW) Hermite spectral expansion, where the weight function does not equal unity. Hermite-based velocity discretizations inherently separate different scales, offering insight into the propagation of free energy (anisotropy or inhomogeneity in the initial distribution function) across scales in both linear and nonlinear processes [43]. Additionally, the AW Hermite expansion, in particular, conserves mass, momentum, and energy (if coupled with a conserving spatial and temporal integrator). The AW Hermite expansion is given by
@@
$\mathcal{H}_{n}(v)$ $\coloneqq(-1)^{n}\exp{\left(v^{2}\right)}\frac{\mathrm{d}^{n}}{\mathrm{d}v^{n}}\exp{\left(v^{2}\right)}.$
The orthogonality and recursive properties of the AW Hermite basis [44, §22] are
@@
with the convention of $\hat{C}_{<0}=0$ and the closure by truncation $\hat{C}_{N_{v}}=0$, such that the closure term introduces an error only in the advection term $v\partial_{x}f$. We later consider other closure ideas in section 3.3. Then, in vector form, we get
$\omega\bar{C}=k\bar{\bar{{A}}}\bar{C}+k\hat{\phi}\bar{B},$
where $\bar{C}\coloneqq[\hat{C}_{0},\hat{C}_{1},\hat{C}_{2},\ldots,\hat{C}_{N_{v}-1}]^{\top}\in\mathbb{R}^{N_{v}},\bar{B}\coloneqq[0\hskip 5.0pt-\frac{1}{\sqrt{2}}\hskip 5.0pt0\ldots 0]^{\top}\in\mathbb{R}^{N_{v}},$ and
@@
Therefore, the Hermite approximation to the response function is
$R^{\mathrm{aw}}_{N_{v}}(\xi)$ $\coloneqq\frac{\sqrt{2}\hat{C}_{0}}{\hat{\phi}}=\frac{\sqrt{2}}{\hat{\phi}}[1\hskip 5.0pt0\hskip 5.0pt0\ldots 0]\left(\xi\mathbb{I}_{N_{v}}-\frac{k}{\sqrt{2}|k|}\bar{\bar{A}}\right)^{-1}\left(\frac{k\hat{\phi}}{\sqrt{2}|k|}\bar{B}\right)$ (15)
@@
The Hermite response function approximations have the correct asymptotic behavior in the fluid limit $|\xi|\gg 1$, i.e. when the phase velocity is much larger than the thermal velocity. In this region, increasing the number of moments improves the order of accuracy of the asymptotic expansion. However, in the adiabatic limit of $|\xi|\ll 1$, the accuracy of the approximation does not improve as we increase the number of Hermite moments $N_{v}$. Additionally, models with even $N_{v}$ are more accurate in the $|\xi|\ll 1$ limit in comparison to models with odd $N_{v}$ (at least to the zeroth order approximation) [45]. The mathematical reasoning is that the eigenvalues of the skew-symmetric advection matrix $ik\bar{\bar{A}}$ defined in Eq. (14) are purely imaginary in the even-dimension case, whereas, in the odd-dimension case, there is an additional unpaired zero eigenvalue. Figure 1 shows the Hermite response function approximation for models with three to six $N_{v}$ moments. The Hermite approximations have purely real poles in the $|\xi|\sim 1$ region that cause the response function to grow unbounded. Furthermore, the kinetic response function $R(\xi)$ has both real and imaginary components, whereas the approximation $R_{N_{v}}^{\mathrm{aw}}(\xi)$ is purely real. A purely real response function indicates that the approximate discrete system cannot sustain a phase shift between the density and electrostatic potential perturbations.
@@ 3.1 Artificial Collisions
where $\nu\in\mathbb{R}_{+}$ is the normalized collisional frequency and $\alpha\in\mathbb{N}_{\geq 1}$ controls the order of dissipation. The artificial collision operator is added to the right-hand side of Eq. (1). If $\alpha\geq 2$, the hypercollisional operator conserves mass, momentum, and energy [51]. The Hermite basis function $\psi_{n}(v)$ defined in Eq. (9) is an eigenfunction of the collisional operator with eigenvalue $-\nu n!/(n-2\alpha+1)!$. The operator in Eq. (16) recovers the LB [29] collisional operator when $\alpha=1$. Similarly,  Camporeale et al. [23] adopted a hypercollisional operator that corresponds to the operator in Eq. (16) with $\alpha=2$ (up to a normalizing factor), which was adopted in various studies, e.g. [15, 16, 11, 12, 13, 17]. Following the recursive Hermite properties in Eq. (11)–(12), see [51], the normalized artificial collisions operator in Hermite space acts as
$\mathcal{C}_{\mathrm{hyper}}(\hat{C}_{n})=-\nu\underbrace{\frac{n!(N_{v}-2\alpha)!}{(n-2\alpha+1)!(N_{v}-1)!}}_{\text{damping rate }\propto n^{2\alpha-1}}\hat{C}_{n}$ (17)
such that the last Hermite moment $\hat{C}_{N_{v}-1}$ damping rate is $\nu\in\mathbb{R}_{+}$. An important feature of the operator is that the damping rate increases with larger $n$ Hermite coefficients, scaling as $n^{2\alpha-1}$.
@@ 3.2 Filtering
Hou and Li [31] use $p=36$ for spatial Fourier smoothing in hydrodynamic simulations, a parameter later adopted for AW Hermite-based solvers (see, e.g., [43, 47, 33, 34, 52]). In this paper, we adopt the same scaling $p=36$. We note that filtering and artificial collisions are fundamentally equivalent methods in the sense that a combination of parameters can be found so that both methods yield very similar damping rates. This is demonstrated in Fig. 2. Specifically, for $N_{v}=20$, artificial collisions with $\alpha=8$ have a damping rate comparable to that of filtering. When $N_{v}=10^{3}$, artificial collisions with $\alpha=18$ show a damping rate similar to that of filtering. In A, we discuss another commonly used filter developed by Klimas [30], which fails to suppress recurrence and recover correct Landau roots unless it is combined with a Fourier-Fourier discretization and certain boundary conditions [32, 4].
@@ 3.3 Hermite Hammett-Perkins Nonlocal Linear Closures
Hammett and Perkins [36] developed a nonlocal linear closure that accounts for collisionless Landau damping effects in a fluid model. This closure approximates the highest moment in the fluid hierarchy by a linear combination of lower-order moments in Fourier space. The transformation between Hermite moments and fluid moments is nonlinear and invertible, such that a linear fluid moment closure corresponds to a nonlinear Hermite moment closure, and vice versa [38]. The transformation from fluid moments to Hermite moments is given by
@@
$\hat{u}$ $=\frac{\hat{C}_{1}}{\hat{C}_{0}}\qquad$ $\Longrightarrow\qquad\hat{C}_{1}=\frac{1}{\sqrt{2}}\hat{u}\hat{n}$
$\hat{p}$ $=\sqrt{2}\left(\sqrt{2}\hat{C}_{2}+\hat{C}_{0}-\frac{\hat{C}_{1}^{2}}{\hat{C}_{0}}\ \right)\qquad$ $\Longrightarrow\qquad\hat{C}_{2}=\frac{\hat{p}}{4}-\frac{\hat{n}}{2}+\frac{\hat{u}^{2}}{2}$
$\hat{q}$ $=\frac{1}{\sqrt{2}}\left(\sqrt{6}\hat{C}_{3}-3\frac{\sqrt{2}\hat{C}_{2}\hat{C}_{1}}{\hat{C}_{0}}+2\frac{\hat{C}_{1}^{3}}{\hat{C}_{0}^{2}}\right)\qquad$ $\Longrightarrow\qquad\hat{C}_{3}=\frac{2\hat{q}}{\sqrt{3}4}+\sqrt{3}\left(\frac{\hat{p}\hat{u}}{4}-\frac{\hat{n}\hat{u}}{2}+2\frac{\hat{u}^{3}}{4}\right)-\frac{4}{\sqrt{3}}\hat{u}^{3}\hat{n}$
@@
Following the work by [38, 24], we generalize the mathematical approach to higher $N_{v}$ dimensions for a nonlocal linear closure of the form
@@
The linear closure coefficients $\mu_{n}\in\mathbb{C}$ are chosen by matching the approximate response function to the analytic response function at the asymptotic limit of $|\xi|\ll 1$, see Eq. (6). The linear closure retains the same order of accuracy at the fluid limit $\xi\gg 1$ as the closure by truncation. Since we solve for the approximate response function symbolically, computational limitations require keeping $N_{m}$ relatively small. To the best of our knowledge, this study is the first to interpret the nonlocal closure approach as a method to mitigate recurrence and suppress filamentation artifacts in kinetic simulations. The concept of nonlocal Hermite closure is similar to the approach proposed by Eliasson [53], which introduced absorbing velocity boundary conditions in Fourier velocity discretizations to reduce recurrence effects.

### v2

@@ 1 Introduction
Kinetic simulations of collisionless plasmas are important for many applications, e.g. astrophysics, geophysics, and laboratory fusion devices. Solving the collisionless Vlasov equation numerically is challenging due to its high dimensionality, Hamiltonian structure, and significant spatial and temporal scale separation. Furthermore, kinetic solvers struggle to accurately capture phase-space mixing, in which the distribution function develops increasingly fine structures in velocity space, known as filamentation [1]. As shown by the solution $f(x,v,t)=f(x-vt,v,0)$ to the free-streaming equation $\partial_{t}f+v\partial_{x}f=0$, these structures cannot be resolved after a finite time determined by the resolution of the velocity space. Filamentation can also cause recurrence in Eulerian solvers (grid or spectral), where the solution is artificially periodic in time [2, 3].
@@
The most common closure for the Hermite spectral approximation is closure by truncation, which assumes that the last spectral moment is zero. However, the Landau root [26] does not emerge as a discrete eigenvalue with the closure by truncation, regardless of velocity resolution. Grant and Feix [1] were the first to show that adding artificial collisions to the finite system with the closure by truncation can restore the correct Landau damping [27, 28, 20]. Previous work by [1, 9] use the Lenard Bernstein (LB) [29] collisional operator. Similarly, recent work by [23, 15, 16, 11, 12, 13, 17] use a hypercollisional operator, which damps more strongly the higher-order Hermite moments in comparison to the LB operator. Filtering techniques, such as Klimas [30] and Hou-Li [31], are also used to suppress recurrence and filamentation artifacts [32, 33, 34, 35, 3]. Another method for incorporating kinetic physics into moment solvers is with a linear nonlocal closure, often referred to as Landau-fluid closures [36, 37]. This approach matches the closure coefficients to the asymptotic limits of the linear kinetic response function. Landau-fluid closures were initially formulated for fluid models incorporating three and four moments and were later generalized to models involving higher moments in Hermite space [38]. Therefore, it becomes essential to incorporate a modified term or closure that introduces dissipation into the discrete system to prevent filamentation issues while ensuring that the chosen approach accurately captures the system’s average dynamics.
@@ 2.2 Linear Response Function
$-i\omega\hat{f}(v)+ikv\hat{f}(v)+ik\hat{\phi}\frac{\mathrm{d}f_{0}}{\mathrm{d}v}=0\qquad\Rightarrow\qquad\hat{f}(v)=k\hat{\phi}\frac{\mathrm{d}f_{0}/\mathrm{d}v}{\omega-kv}.$ (4)
For a Maxwellian equilibrium, that is, $f_{0}(v)=\exp\left(-v^{2}/2\right)/\sqrt{2\pi}$, the perturbed density is described via the following linear kinetic response function $R(\xi)$:
@@
where the plasma dispersion function is $Z(\xi)=\pi^{-1/2}\int_{\mathbb{R}}\exp(-s^{2})/(s-\xi)\mathrm{d}s$ for $\mathrm{Im}(\xi)>0$ and is analytically continued for $\mathrm{Im}(\xi)\leq 0$. There is a subtle distinction between defining $\xi\coloneqq\omega/(\sqrt{2}k)$ and $\xi\coloneqq\omega/(\sqrt{2}|k|)$. We adopt the latter definition because it allows the use of the plasma dispersion function $Z(\xi)$ as originally defined by Fried and Conte [42]. In contrast, the former definition requires a redefinition of $Z(\xi)$; see Hunana et al. [40, §2.2]. The asymptotic expansion of the response function is given by [40, §3]:
@@ 2.3 Asymmetrically Weighted Hermite Expansion
We discretize the electron distribution function in velocity space via an asymmetrically weighted (AW) Hermite spectral expansion, where the weight function does not equal unity. Hermite-based velocity discretizations inherently separate different scales (analogous to other orthogonal bases, e.g. Fourier), offering insight into the propagation of free energy (anisotropy or inhomogeneity in the initial distribution function) across scales in both linear and nonlinear processes [43]. Additionally, the AW Hermite expansion, in particular, conserves mass, momentum, and energy (if coupled with a conserving spatial and temporal integrator). The AW Hermite expansion is given by
@@
$\mathcal{H}_{n}(v)$ $\coloneqq(-1)^{n}\exp{\left(v^{2}\right)}\frac{\mathrm{d}^{n}}{\mathrm{d}v^{n}}\exp{\left(-v^{2}\right)}.$
The limit relation of the AW Hermite basis function [44, §22.15] is given by
$\lim_{n\to\infty}\psi_{n}(v)\propto\begin{cases}\cos(\sqrt{n}v)\exp\left(-\frac{v^{2}}{2}\right)\qquad&\text{if n is even},\\ \sin(\sqrt{n-1}v)\exp\left(-\frac{v^{2}}{2}\right)\qquad&\text{if n is odd},\end{cases}$ (10)
such that the shortest wave-length $\lambda_{v}$ in velocity space is resolved by the last Hermite basis of $N_{v}-1$, i.e $\lambda_{v}\approx 2\pi/\sqrt{N_{v}-1}$. The orthogonality and recursive properties of the AW Hermite basis [44, §22.2–22.8] are
@@
with the convention of $\hat{C}_{<0}=0$ and the closure by truncation $\hat{C}_{N_{v}}=0$, such that the closure term introduces an error only in the advection term $v\partial_{x}f$. We later consider other closure ideas in section 3.3. Then, in vector form, Eq. (14) becomes
$\omega\bar{C}=k\bar{\bar{{A}}}\bar{C}+k\hat{\phi}\bar{B}\qquad\Rightarrow\qquad\left(\xi\mathbb{I}_{N_{v}}-\frac{k}{\sqrt{2}|k|}\bar{\bar{A}}\right)\bar{C}=\frac{k\hat{\phi}}{\sqrt{2}|k|}\bar{B},$ (15)
where $\bar{C}\coloneqq[\hat{C}_{0},\hat{C}_{1},\hat{C}_{2},\ldots,\hat{C}_{N_{v}-1}]^{\top}\in\mathbb{R}^{N_{v}},\bar{B}\coloneqq[0\hskip 5.0pt-\frac{1}{\sqrt{2}}\hskip 5.0pt0\ldots 0]^{\top}\in\mathbb{R}^{N_{v}},\mathbb{I}_{N_{v}}\in\mathbb{R}^{N_{v}\times N_{v}}$ is the identity matrix, the variable $\xi$ is defined in Eq. (5), and
@@
Therefore, inserting Eq. (15) in the response function definition in Eq. (5) results in the following Hermite approximation to the response function
$R^{\mathrm{aw}}_{N_{v}}(\xi)$ $\coloneqq\frac{\hat{n}}{\hat{\phi}}=\frac{\sqrt{2}\hat{C}_{0}}{\hat{\phi}}=\frac{\sqrt{2}}{\hat{\phi}}[1\hskip 5.0pt0\hskip 5.0pt0\ldots 0]\left(\xi\mathbb{I}_{N_{v}}-\frac{k}{\sqrt{2}|k|}\bar{\bar{A}}\right)^{-1}\left(\frac{k\hat{\phi}}{\sqrt{2}|k|}\bar{B}\right)$ (17)
@@
The Hermite response function approximations have the correct asymptotic behavior in the fluid limit $|\xi|\gg 1$, i.e. when the phase velocity is much larger than the thermal velocity. In this region, increasing the number of moments improves the order of accuracy of the asymptotic expansion. However, in the adiabatic limit of $|\xi|\ll 1$, the accuracy of the approximation does not improve as we increase the number of Hermite moments $N_{v}$. Additionally, models with even $N_{v}$ are more accurate in the $|\xi|\ll 1$ limit in comparison to models with odd $N_{v}$ (at least to the zeroth order approximation) [45]. The mathematical reasoning is that the eigenvalues of the skew-symmetric advection matrix $ik\bar{\bar{A}}$ defined in Eq. (16) are purely imaginary in the even-dimension case, whereas, in the odd-dimension case, there is an additional unpaired zero eigenvalue. Figure 1 shows the Hermite response function approximation for models with three to six moments. The Hermite approximations have purely real poles in the $|\xi|\sim 1$ region that cause the response function to grow unbounded. Furthermore, the kinetic response function $R(\xi)$ has both real and imaginary components for purely real $\xi$, whereas the approximation $R_{N_{v}}^{\mathrm{aw}}(\xi)$ is purely real for purely real $\xi$. A purely real response function indicates that the approximate discrete system cannot sustain a phase shift between the density and electrostatic potential perturbations.
@@ 3.1 Artificial Collisions
where $\nu\in\mathbb{R}_{+}$ is the normalized collisional frequency and $\alpha\in\mathbb{N}_{\geq 1}$ controls the order of dissipation. The artificial collision operator is added to the right-hand side of Eq. (1). If $\alpha\geq 2$, the hypercollisional operator conserves mass, momentum, and energy [51]. The Hermite basis function $\psi_{n}(v)$ defined in Eq. (9) is an eigenfunction of the collisional operator with eigenvalue $-\nu n!/(n-2\alpha+1)!$. The operator in Eq. (18) recovers the LB [29] collisional operator when $\alpha=1$. Similarly,  Camporeale et al. [23] adopted a hypercollisional operator that corresponds to the operator in Eq. (18) with $\alpha=2$ (up to a normalizing factor), which was adopted in various studies, e.g. [15, 16, 11, 12, 13, 17]. Following the recursive Hermite properties in Eq. (12)–(13), see Funaro and Manzini [51], the normalized artificial collisions operator in Hermite space acts as
$\mathcal{C}_{\mathrm{hyper}}(\hat{C}_{n})=-\nu\underbrace{\frac{n!}{(n-2\alpha+1)!}\underbrace{\frac{(N_{v}-2\alpha)!}{(N_{v}-1)!}}_{\text{normalizing factor}}}_{\text{damping rate }\propto n^{2\alpha-1}}\hat{C}_{n}$ (19)
such that the last Hermite moment $\hat{C}_{N_{v}-1}$ damping rate is $\nu\in\mathbb{R}_{+}$. The hypercollisional operator should always be applied in a convergence sense, ensuring that it does not substantially alter the collisionless physics of interest. An important feature of the operator is that the damping rate increases with larger $n$ Hermite coefficients, scaling as $n^{2\alpha-1}$.
@@ 3.2 Filtering
Hou and Li [31] use $p=36$ for spatial Fourier smoothing in hydrodynamic simulations, a parameter later adopted for AW Hermite-based solvers (see, e.g., [43, 47, 33, 34, 52]). In this paper, we adopt the same scaling $p=36$. We note that filtering and artificial collisions are fundamentally equivalent methods in the sense that a combination of parameters can be found so that both methods yield very similar damping rates. This is demonstrated in Fig. 2. Specifically, for $N_{v}=20$, artificial collisions with $\alpha=8$ have a damping rate comparable to that of filtering. When $N_{v}=10^{3}$, artificial collisions with $\alpha=18$ show a damping rate similar to that of filtering. In A, we discuss another commonly used filter developed by Klimas [30]. The Klimas filter does not correspond to a true velocity space dissipation and hence the Landau root is not a discrete mode of the Klimas filtered AW Hermite discretization with closure by truncation.
@@ 3.3 Hermite Hammett-Perkins Nonlocal Linear Closures
Hammett and Perkins [36] developed a nonlocal linear closure that accounts for collisionless Landau damping effects in a fluid model. This closure approximates the highest moment in the fluid hierarchy by a linear combination of lower-order moments in Fourier space. The transformation between Hermite moments and fluid moments is nonlinear and invertible, such that a linear fluid moment closure corresponds to a nonlinear Hermite moment closure, and vice versa [38, §3]. The transformation from fluid moments to Hermite moments is
@@
$\hat{u}$ $=\frac{\hat{C}_{1}}{\hat{C}_{0}}\qquad$ $\Longrightarrow\qquad\hat{C}_{1}=\frac{\hat{u}\hat{n}}{\sqrt{2}}$
$\hat{p}$ $=2\hat{C}_{2}+\sqrt{2}\hat{C}_{0}-\frac{\sqrt{2}\hat{C}_{1}^{2}}{\hat{C}_{0}}\qquad$ $\Longrightarrow\qquad\hat{C}_{2}=\frac{\hat{p}-\hat{n}+\hat{u}^{2}\hat{n}}{2}$
$\hat{q}$ $=2\sqrt{3}\hat{C}_{3}-\frac{6\hat{C}_{2}\hat{C}_{1}}{\hat{C}_{0}}+\frac{2\sqrt{2}\hat{C}_{1}^{3}}{\hat{C}_{0}^{2}}\qquad$ $\Longrightarrow\qquad\hat{C}_{3}=\frac{\hat{q}+3\hat{u}\left(\hat{p}-\hat{n}\right)+\hat{u}^{3}\hat{n}}{2\sqrt{3}}$
@@
Following the work by Smith [38], Hammett et al. [24], we generalize the mathematical approach to higher $N_{v}$ dimensions for a nonlocal linear closure of the form
@@
The number of lower order Hermite moments used to close the equations, which corresponds to the number of adjustable parameters in the linear closure, is denoted by $N_{m}\in\mathbb{N}_{>0}$. The linear closure coefficients $\mu_{n}\in\mathbb{C}$ are chosen by matching the approximate response function to the analytic response function at the asymptotic limit of $|\xi|\ll 1$, see Eq. (6). The linear closure retains the same order of accuracy at the fluid limit $\xi\gg 1$ as the closure by truncation. Since we solve for the approximate response function symbolically, computational limitations require keeping $N_{m}$ relatively small. To the best of our knowledge, this study is the first to interpret the nonlocal closure approach as a method to mitigate recurrence and suppress filamentation artifacts in kinetic simulations. The concept of nonlocal Hermite closure is similar to the approach proposed by Eliasson [53], which introduced absorbing velocity boundary conditions in Fourier velocity discretizations to reduce recurrence effects.